Hello,
I am not really sure where to put this question...what topic it would fall under...
Here is my problem:
-3 1/2x + 2 3/4 + 1/2x = 3 7/8
I tried several ways to solve this but I could not come up with the correct answer... and I am not sure if I am approaching the problem correctly.
This is from an algebra one course...
Thank you very much,
Diane M.

Found 2 solutions by nerdybill, edjones:
Answer by nerdybill(2707)

Your equation is ambiguous. People reading it can interpret it in potentially more than one way. Correct me if I interpreted your problem wrong.
.
Starting with:
-3 1/2x + 2 3/4 + 1/2x = 3 7/8
Your first step is to change the mixed numbers into improper fractions:
-3 1/2 = -(6/2 + 1/2) = -7/2
2 3/4 = 8/4 + 3/4 = 11/4
3 7/8 = 24/8 + 7/8 = 31/8
.
Now, we can rewrite:
-3 1/2x + 2 3/4 + 1/2x = 3 7/8
As
(-7/2)x + 11/4 + (1/2)x = 31/8
Multiply both sides by 8 to get rid of all denominators:
8[(-7/2)x + 11/4 + (1/2)x] = 8[31/8]
-28x + 22 + 4x = 31
-24x + 22 = 31
22 = 24x + 31
-9 = 24x
-9/24 = x
-3/8 = x

Answer by edjones(4080)

(-3 1/2)x + 2 3/4 + (1/2)x = 3 7/8 I assume that this is the problem.
-7x/2 + 11/4 + x/2 = 31/8
-28x+22+4x=31 multiply each side by 8 to eliminate fractions.
-24x=9
x=-9/24
=-3/8
.
Ed